# Qubit encoding for a mixture of localized functions

### v1

@@ II.3.1 Discrete Slater functions
The discrete SFs for five qubits having various decay rates are plotted in Fig. 4(b) as examples.
@@ II.3.2 Discrete Lorentzian functions
$a$ is a real positive parameter. This is a period-$N$ function having peaks at the origin and its duplicated points, as well as the discrete SF. The discrete LFs for five qubits having various decay rates are plotted in Fig. 4(c) as examples.
@@
generates the LF state from initialized qubits. $\mathrm{QFT}$ on the right-hand side of Eq. (23) can be replaced by $\mathrm{QFT}^{\dagger}.$ As seen in Figs. 4(b) and (c), a discrete SF having a larger decay rate corresponds to a discrete LF having a wider peak. In fact, from Eq. (21), the discrete LF behaves near the origin $(|j|\ll N)$ approximately as $L_{j}(n_{q},a)\propto 1/(j^{2}+\gamma_{L}(a)^{2}),$ where $\gamma_{L}(a)\equiv(N/\pi)\sinh(a/2)$ is a monotonically increasing function of $a.$ In what follows, we denote $a$ for $L(n_{q},a)$ simply as the decay rate of the discrete LF since no confusion would occur. The overlap between two LFs is given as a simple expression, as derived in Appendix G.2.
The circuit depth of $U^{(\mathrm{L})}$ implemented as Eq. (23) is, from Eq. (20),
$\mathrm{depth}(U^{(\mathrm{L})})=\mathcal{O}(n_{q}\log n_{q}).$ (24)
@@ II.4.1 Probabilistic encoding
We know the unitaries $U^{(\mathrm{S})}$ and $U^{(\mathrm{L})}$ for generating an SF and an LF at the origin, respectively, as described in Sect. II.3. We can thus construct explicitly the circuit for encoding probabilistically an LC $|\psi_{\mathrm{lc}}\rangle$ of SFs and/or LFs according to Sect. II.2. Although the circuit of $U^{(\mathrm{S})}$ is shallower than that of $U^{(\mathrm{L})}$ [see Eqs. (20) and (24)], the circuits for displaced single SF and LF have the common scaling $\mathcal{O}(n_{q}\log n_{q})$ due to the QFT operations, as found in Fig. 3(a). When we adopt the LFs as the basis functions for the expansion in Eq. (7), $\mathrm{QFT}^{\dagger}$ in Fig. 3(a) cancels $\mathrm{QFT}$ in $U^{(\mathrm{L})},$ as understood from Eq. (23). This means that we can generate the displaced single LF by implementing $U^{(\mathrm{S})}$ and performing $\mathrm{QFT}$ only once. This favorable property of the LFs are inherited to the circuit $\mathcal{C}_{\mathrm{lc}}$ in Fig. 3(b) for the desired LC. We therefore adopt the LFs as the basis set in what follows:
@@
Let us estimate the scaling of depth of $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ with respect to $n_{q},n_{\mathrm{loc}},$ and $n_{\mathrm{A}}.$ The implementation technique for MCM1 gates is also applicable to the part for the phase shift gates. The controlled $y$ rotations among the ancillae can be implemented with a depth of $\mathcal{O}(n_{\mathrm{loc}}n_{\mathrm{A}})$ [45, 46]. The total depth is thus estimated to be
$\mathrm{depth}\left(\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}\right)$ $=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}\log n_{\mathrm{loc}},n_{\mathrm{loc}}\log n_{q},n_{q}\log n_{q}\right)\right),$ (26)
@@
$=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}^{2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{2}\log n_{q},n_{q}\log n_{q}\right)\right)$ (27)
@@ II.4.2 Deterministic encoding
Since the QFT gate in $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L,AR})}$ does not act on the $n_{\mathrm{A}}+1$ ancillae, the amplification operator $Q$ for QAA [12, 13] can be constructed from the partial circuit $\mathcal{U}^{(\mathrm{L,AR})}$ [see Fig. 6(a)] as
$Q\equiv-\mathcal{U}^{(\mathrm{L,AR})}S_{0}\mathcal{U}^{(\mathrm{L,AR})\dagger}S_{0},$ (28)
where
@@
The encoding is still probabilistic for $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L},m)}$ with a generic combination of $m$ and $\theta_{\mathrm{AR}}.$ We can determinize it, however, by choosing an appropriate combination of them. Specifically, the optimal value $m_{\mathrm{opt}}$ of $m$ is found from the weight $w$ of success state in $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ according to Eq. (4). With this, we find $\theta_{\mathrm{AR,opt}}$ from Eq. (5) for $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L},m_{\mathrm{opt}})}$ to obtain the deterministic circuit $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L,det})},$ which is the main result of this study.
Let us estimate the scaling of depth of $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L},m)}$ for a given $m.$ From $\mathrm{depth}(\mathcal{U}^{(\mathrm{L,AR})})=\mathrm{depth}(\mathcal{U}^{(\mathrm{L})})$ [see Fig. 5(b)] and $\mathrm{depth}(S_{0})=\mathcal{O}(n_{\mathrm{A}})$ [45, 46], we find $\mathrm{depth}(Q)=\mathcal{O}(n_{\mathrm{loc}}\max(n_{\mathrm{A}},\log n_{q})).$ The total depth thus scales as
@@
$=\mathcal{O}\left(\max(mn_{\mathrm{loc}}\log n_{\mathrm{loc}},mn_{\mathrm{loc}}\log n_{q},n_{q}\log n_{q})\right).$ (30)
@@
$=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}^{3/2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{3/2}\log n_{q},n_{q}\log n_{q}\right)\right)$ (31)
@@ II.5 Resource estimation for quantum chemistry in real space
One of the intriguing applications of our scheme is the encoding of a one-electron orbital in a molecular system for quantum chemistry in real space [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. As explained in Appendix E, our scheme is straightforwardly extended to a three-dimensional case by employing product basis functions. Before we start a simulation of real-time dynamics or an energy minimization procedure toward the ground state of a target molecule, we have to prepare an initial many-electron state. In the first-quantized formalism for a quantum computer, explicit antisymmetrization of an initial state is mandatory [25, 26, 24]. The efficient antisymmetrization scheme proposed by Berry et al. [26] assumes that all the occupied MOs have been prepared. If we adopt such a scheme, techniques for generating MOs of various shapes have to be established. In this subsection, we analyze the computational cost that is required for encoding MOs as LCs of LFs.
@@ II.5.1 Probabilistic encoding
As discussed in Ref. [18], the required number of qubits for encoding a one-electron orbital scales typically as $n_{qe}=\mathcal{O}(\log(n_{\mathrm{el}}^{1/3}/\Delta x))$ for each direction. $n_{\mathrm{el}}$ is the number of electrons contained in a target molecule. The data register thus requires $n_{q}=3n_{qe}$ qubits.
@@
$=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}\log n_{\mathrm{loc}},n_{\mathrm{loc}}\log\log n_{\mathrm{el}},\left(\log n_{\mathrm{el}}\right)\log\log n_{\mathrm{el}}\right)\right).$ (32)
For a case where we want to encode an MO, which is a chemically motivated LC of one-electron atomic-like orbitals, we can estimate the computational cost by considering only the electron number. Specifically, if the desired MO is delocalized over the molecule such as the $\pi$ network in a hydrocarbon molecule, it is made up of the contributions from the constituent atoms with roughly equal weights, as exemplified in Fig. 7(a). The number of basis functions for it thus scales the same way as the molecule size: $n_{\mathrm{loc}}=\mathcal{O}(n_{\mathrm{el}}).$ The circuit depth in this case is, from Eq. (32),
$\mathrm{depth}\left(\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}\ \mathrm{for\ a\ delocalized\ MO}\right)$
$=\mathcal{O}\left(n_{\mathrm{el}}\log n_{\mathrm{el}}\right).$ (33)
If the desired MO is localized, on the other hand, such as one in a transition metal complex depicted in Fig. 7(b), we have clearly $n_{\mathrm{loc}}=\mathcal{O}(1).$ Equation (32) in this case reads
$\mathrm{depth}\left(\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}\ \mathrm{for\ a\ localized\ MO}\right)$
$=\mathcal{O}\left(\left(\log n_{\mathrm{el}}\right)\log\log n_{\mathrm{el}}\right).$ (34)
As for the computational time of the probabilistic encoding, we find from Eq. (27)
@@
$=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}^{2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{2}\log\log n_{\mathrm{el}},(\log n_{\mathrm{el}})\log\log n_{\mathrm{el}}\right)\right)$ (35)
@@
Generic $\mathcal{O}(\mathrm{max}(n_{\mathrm{loc}}\log n_{\mathrm{loc}},n_{\mathrm{loc}}\log\log n_{\mathrm{el}},(\log n_{\mathrm{el}})\log\log n_{\mathrm{el}}))$ $\mathcal{O}(\mathrm{max}(n_{\mathrm{loc}}^{2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{2}\log\log n_{\mathrm{el}},(\log n_{\mathrm{el}})\log\log n_{\mathrm{el}}))$
@@
Localized $\mathcal{O}((\log n_{\mathrm{el}})\log\log n_{\mathrm{el}})$ $\mathcal{O}((\log n_{\mathrm{el}})\log\log n_{\mathrm{el}})$
@@
Generic $\mathcal{O}(\mathrm{max}(n_{\mathrm{loc}}^{3/2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{3/2}\log\log n_{\mathrm{el}},(\log n_{\mathrm{el}})\log\log n_{\mathrm{el}}))$
@@
Localized $\mathcal{O}((\log n_{\mathrm{el}})\log\log n_{\mathrm{el}})$
@@ II.6 Classical computation of finding the optimal LC for a target function
where $\bm{d},\bm{a},$ and $\bm{k}_{\mathrm{c}}$ represent collectively the expansion coefficients, decay rates, and centers, respectively, of $n_{\mathrm{loc}}$ LFs. For a fixed $n_{\mathrm{loc}}$, we want to find on a classical computer the combination of these parameters so that the squared overlap
$F(\bm{d},\bm{a},\bm{k}_{\mathrm{c}})$ $\equiv|\langle\psi[\bm{d},\bm{a},\bm{k}_{\mathrm{c}}]|\psi_{\mathrm{ideal}}\rangle|^{2}$ (37)
between the trial state and the target state is maximized with obeying the normalization condition $\||\psi[\bm{d},\bm{a},\bm{k}_{\mathrm{c}}]\rangle\|^{2}=1.$
There may exist various numerical procedures for finding the optimal parameters. Here we provide an outline of a possible one that exploits the fact that the objective function $F$ is in a quadratic form with respect to $\bm{d}.$ Specifically, from $g_{\ell}\equiv\sum_{j=0}^{N-1}\psi_{\mathrm{ideal}}(x_{j})L_{j}(a_{\ell},n_{q})$ for each $\ell$ and the overlaps between the LF states (see Appendix G.2), we construct the matrices $G$ and $S$ as $G_{\ell\ell^{\prime}}\equiv g_{\ell}g_{\ell^{\prime}}$ and $S_{\ell\ell^{\prime}}\equiv V(a_{\ell},a_{\ell^{\prime}},k_{\mathrm{c}\ell}-k_{\mathrm{c}\ell^{\prime}}),$ respectively. The stationarity conditions for $F$ with respect to the expansion coefficients lead to the generalized eigenvalue problem:
$G\bm{c}=\lambda S\bm{c},$ (38)
where $\lambda$ is the Lagrange multiplier responsible for the normalization condition and $\bm{c}$ is the solution of this equation. The solution out of $n_{\mathrm{loc}}$ that maximizes $F$ gives the globally optimal coefficients for fixed $\bm{a}$ and $\bm{k}_{\mathrm{c}}:$
@@
As an example for the procedure described above, we tried to find the optimal LC to express the sum of two Gaussian functions:
$\psi_{\mathrm{ideal}}(x_{j})\propto\exp\left(-\frac{(j-16)^{2}}{9}\right)$
$+0.4\exp\left(-\frac{(j-8)^{2}}{4}\right)\ (j=0,\dots,31)$ (40)
for $n_{q}=5$ by using $n_{\mathrm{loc}}=3$ LFs. We obtained the decay rates $a_{0}=0.360,a_{1}=0.490,$ and $a_{2}=1.672$ and the centers $k_{\mathrm{c}0}=8,k_{\mathrm{c}1}=16,$ and $k_{\mathrm{c}2}=12$ for the LFs. Their coefficients were $d_{0}=0.417,d_{1}=1.23,$ and $d_{2}=-0.507.$ The target state $|\psi_{\mathrm{ideal}}\rangle$ and the optimal state $|\psi[\bm{d},\bm{a},\bm{k}_{\mathrm{c}}]\rangle$ are plotted in Fig. 8. The squared overlap between the target and optimal states was found to be $0.992$
Figure 8: Target state $|\psi_{\mathrm{ideal}}\rangle$ for $n_{q}=5$ in Eq. (40) and the LC of LF states ($n_{\mathrm{loc}}=3$) $|\psi[\bm{d},\bm{a},\bm{k}_{\mathrm{c}}]\rangle$ that optimally approximates the target state. The contribution from each LF state is also shown.

### v2

@@ II.3.1 Discrete Slater functions
The discrete SFs for five data qubits having various decay rates are plotted in Fig. 3(b) as examples.
@@ II.3.2 Discrete Lorentzian functions
$a$ is a real positive parameter. This period-$N$ function has peaks at the origin and its duplicated points, as well as the discrete SF. The discrete LFs for five data qubits having various decay rates are plotted in Fig. 3(c) as examples.
@@
generates the LF state from initialized qubits. $\mathrm{QFT}$ on the right-hand side of Eq. (21) can be replaced by $\mathrm{QFT}^{\dagger}.$ As seen in Figs. 3(b) and (c), a discrete SF having a larger decay rate corresponds to a discrete LF having a wider peak. In fact, from Eq. (19), the discrete LF behaves near the origin $(|j|\ll N)$ approximately as $L_{j}(n_{q},a)\propto 1/(j^{2}+\gamma_{L}(a)^{2}),$ where $\gamma_{L}(a)\equiv(N/\pi)\sinh(a/2)$ is a monotonically increasing function of $a.$ In what follows, we denote $a$ for $L(n_{q},a)$ simply as the decay rate of the discrete LF since no confusion would occur. The overlap between two LFs is given as a simple expression, as derived in Appendix G.2. The circuit depth of $U^{(\mathrm{L})}$ implemented as Eq. (21) is, from Eq. (18),
$\mathrm{depth}(U^{(\mathrm{L})})=\mathcal{O}(n_{q}).$ (22)
@@ II.4.1 Probabilistic encoding
We know the unitaries $U^{(\mathrm{S})}$ and $U^{(\mathrm{L})}$ for generating an SF and an LF at the origin, respectively, as described in Sect. II.3. We can thus construct explicitly the circuit for encoding probabilistically an LC $|\psi_{\mathrm{lc}}\rangle$ of SFs and/or LFs according to Sect. II.2. Although the circuit of $U^{(\mathrm{S})}$ is shallower than that of $U^{(\mathrm{L})}$ [see Eqs. (18) and (22)], the circuits for displaced single SF and LF have the common scaling $\mathcal{O}(n_{q})$ due to the QFT operations, as found in Fig. 2(a). When we adopt the LFs as the basis functions for the expansion in Eq. (7), $\mathrm{QFT}^{\dagger}$ in Fig. 2(a) cancels $\mathrm{QFT}$ in $U^{(\mathrm{L})},$ as understood from Eq. (21). This means that we can generate the displaced single LF by implementing $U^{(\mathrm{S})}$ and performing $\mathrm{QFT}$ only once. This favorable property of the LFs is inherited by the circuit $\mathcal{C}_{\mathrm{lc}}$ in Fig. 2(b) for the desired LC. We therefore adopt the LFs as the basis set in what follows:
@@
Let us estimate the scaling of depth of $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ with respect to $n_{q},n_{\mathrm{loc}},$ and $n_{\mathrm{A}}.$ The implementation technique for MCM1 gates is also applicable to the part for the phase shift gates. The controlled $y$ rotations among the ancillae can be implemented with a depth of $\mathcal{O}(n_{\mathrm{loc}}n_{\mathrm{A}})$ [48, 49, 50]. The total depth is thus estimated to be
$\mathrm{depth}\left(\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}\right)$ $=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}\log n_{\mathrm{loc}},n_{\mathrm{loc}}\log n_{q},n_{q}\right)\right),$ (24)
@@
$=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}^{2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{2}\log n_{q},n_{q}\right)\right).$ (25)
@@ II.4.2 Deterministic encoding
Since the QFT gate in $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L,AR})}$ does not act on the $n_{\mathrm{A}}+1$ ancillae, the amplification operator $Q$ for QAA [12, 13] can be constructed from the partial circuit $\mathcal{U}^{(\mathrm{L,AR})}$ [see Fig. 5(a)] as $Q\equiv-\mathcal{U}^{(\mathrm{L,AR})}S_{0}\mathcal{U}^{(\mathrm{L,AR})\dagger}S_{0},$ where
@@
The encoding is still probabilistic for $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L},m)}$ with a generic combination of $m$ and $\theta_{\mathrm{AR}}.$ We can determinize it, however, by choosing an appropriate combination of them. Specifically, the optimal value $m_{\mathrm{opt}}$ of $m$ is found from the weight $w$ of success state in $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ according to Eq. (4). With this, we find $\theta_{\mathrm{AR,opt}}$ from Eq. (5) for $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L},m_{\mathrm{opt}})}$ to obtain the deterministic circuit $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L,det})},$ which is the main result of this study. We would like to emphasize here that we are able to evaluate $w$ analytically without performing QAE in this case since we know the expressions for the circuit parameters in $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}$ (see Appendix A in Ref. [11]).
Let us estimate the scaling of the depth of $\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L},m)}$ for a given $m.$ From $\mathrm{depth}(\mathcal{U}^{(\mathrm{L,AR})})=\mathrm{depth}(\mathcal{U}^{(\mathrm{L})})$ [see Fig. 4(b)] and $\mathrm{depth}(S_{0})=\mathcal{O}(n_{\mathrm{A}})$ [48, 49, 50], we find $\mathrm{depth}(Q)=\mathcal{O}(n_{\mathrm{loc}}\max(n_{\mathrm{A}},\log n_{q})).$ The total depth thus scales as
@@
$=\mathcal{O}\left(\max(mn_{\mathrm{loc}}\log n_{\mathrm{loc}},mn_{\mathrm{loc}}\log n_{q},n_{q})\right).$ (27)
@@
$=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}^{3/2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{3/2}\log n_{q},n_{q}\right)\right)$ (28)
@@ II.5 Resource estimation for quantum chemistry in real space
One of the intriguing applications of our scheme is the encoding of a one-electron spatial orbital in a molecular system for quantum chemistry in real space [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. As explained in Appendix E, our scheme is straightforwardly extended to a three-dimensional case by employing product basis functions. Before we start a simulation of real-time dynamics or an energy minimization procedure toward the ground state of a target molecule, we have to prepare an initial many-electron state. In the first-quantized formalism for a quantum computer, explicit antisymmetrization of an initial state is mandatory [25, 26, 24]. The efficient antisymmetrization scheme proposed by Berry et al. [26] assumes that all the occupied MOs have been prepared. If we adopt such a scheme, techniques for generating MOs of various shapes have to be established. In this subsection, we analyze the computational cost that is required for encoding MOs as LCs of LFs.
@@ II.5.1 Probabilistic encoding
As discussed in Ref. [18], the required number of qubits for encoding a one-electron spatial orbital scales typically as $n_{qe}=\mathcal{O}(\log(n_{\mathrm{el}}^{1/3}/\Delta x))$ for each direction. $n_{\mathrm{el}}$ is the number of electrons contained in a target molecule. The data register thus requires $n_{q}=3n_{qe}$ qubits.
@@
$=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}\log n_{\mathrm{loc}},n_{\mathrm{loc}}\log\log n_{\mathrm{el}},\log n_{\mathrm{el}}\right)\right).$ (29)
For a case where we want to encode an MO, which is a chemically motivated LC of one-electron atomic-like orbitals, we can estimate the computational cost by considering only the electron number. Specifically, if the desired MO is delocalized over the molecule such as the $\pi$ network in a hydrocarbon molecule, it is made up of the contributions from the constituent atoms with roughly equal weights, as exemplified in Fig. 6(a). The number of basis functions for it thus scales the same way as the molecule size: $n_{\mathrm{loc}}=\mathcal{O}(n_{\mathrm{el}}).$ The circuit depth in this case is, from Eq. (29), $\mathrm{depth}(\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}\ \mathrm{for\ a\ delocalized\ MO})=\mathcal{O}(n_{\mathrm{el}}\log n_{\mathrm{el}}).$ If the desired MO is localized, on the other hand, such as one in a transition metal complex depicted in Fig. 6(b), we have clearly $n_{\mathrm{loc}}=\mathcal{O}(1).$ Equation (29) in this case reads $\mathrm{depth}(\mathcal{C}_{\mathrm{lc}}^{(\mathrm{L})}\ \mathrm{for\ a\ localized\ MO})=\mathcal{O}(\log n_{\mathrm{el}}).$
As for the computational time of the probabilistic encoding, we find from Eq. (25),
@@
$=\mathcal{O}\left(\max\left(n_{\mathrm{loc}}^{2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{2}\log\log n_{\mathrm{el}},\log n_{\mathrm{el}}\right)\right).$ (30)
@@
Generic $\mathcal{O}(\mathrm{max}(n_{\mathrm{loc}}\log n_{\mathrm{loc}},n_{\mathrm{loc}}\log\log n_{\mathrm{el}},\log n_{\mathrm{el}}))$ $\mathcal{O}(\mathrm{max}(n_{\mathrm{loc}}^{2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{2}\log\log n_{\mathrm{el}},\log n_{\mathrm{el}}))$
@@
Localized $\mathcal{O}(\log n_{\mathrm{el}})$ $\mathcal{O}(\log n_{\mathrm{el}})$
@@
Generic $\mathcal{O}(\mathrm{max}(n_{\mathrm{loc}}^{3/2}\log n_{\mathrm{loc}},n_{\mathrm{loc}}^{3/2}\log\log n_{\mathrm{el}},\log n_{\mathrm{el}}))$
@@
Localized $\mathcal{O}(\log n_{\mathrm{el}})$
@@ II.6 Classical computation of finding the optimal LC for a target function
where $\bm{d},\bm{a},$ and $\bm{k}_{\mathrm{c}}$ represent the expansion coefficients, decay rates, and centers, respectively, of $n_{\mathrm{loc}}$ LFs. For a fixed $n_{\mathrm{loc}}$, we want to find on a classical computer the combination of these parameters so that the squared overlap $F(\bm{d},\bm{a},\bm{k}_{\mathrm{c}})\equiv|\langle\psi[\bm{d},\bm{a},\bm{k}_{\mathrm{c}}]|\psi_{\mathrm{ideal}}\rangle|^{2}$ between the trial state and the target state is maximized with obeying the normalization condition $\||\psi[\bm{d},\bm{a},\bm{k}_{\mathrm{c}}]\rangle\|^{2}=1.$
There may exist various numerical procedures for finding the optimal parameters. Here we provide an outline of a possible one that exploits the fact that the objective function $F$ is in a quadratic form with respect to $\bm{d}.$ Specifically, from $g_{\ell}\equiv\sum_{j=0}^{N-1}\psi_{\mathrm{ideal}}(x_{j})L_{j}(a_{\ell},n_{q})$ for each $\ell$ and the overlaps between the LF states (see Appendix G.2), we construct the matrices $G$ and $S$ as $G_{\ell\ell^{\prime}}\equiv g_{\ell}g_{\ell^{\prime}}$ and $S_{\ell\ell^{\prime}}\equiv V(a_{\ell},a_{\ell^{\prime}},k_{\mathrm{c}\ell}-k_{\mathrm{c}\ell^{\prime}}),$ respectively. The stationarity conditions for $F$ with respect to the expansion coefficients lead to the generalized eigenvalue problem: $G\bm{c}=\lambda S\bm{c},$ where $\lambda$ is the Lagrange multiplier responsible for the normalization condition and $\bm{c}$ is the solution of this equation. The solution out of $n_{\mathrm{loc}}$ that maximizes $F$ gives the globally optimal coefficients for fixed $\bm{a}$ and $\bm{k}_{\mathrm{c}}:$
@@
As an example for the procedure described above, we tried to find the optimal LC to express the sum of two Gaussian functions: $\psi_{\mathrm{ideal}}(x_{j})\propto\exp(-(j-16)^{2}/9)+0.4\exp(-(j-8)^{2}/4)$ at $j=0,\dots,31$ for $n_{q}=5$ data qubits by using $n_{\mathrm{loc}}=3$ LFs. We obtained the decay rates $a_{0}=0.360,a_{1}=0.490,$ and $a_{2}=1.672$ and the centers $k_{\mathrm{c}0}=8,k_{\mathrm{c}1}=16,$ and $k_{\mathrm{c}2}=12$ for the LFs. Their coefficients were $d_{0}=0.417,d_{1}=1.23,$ and $d_{2}=-0.507.$ The target state $|\psi_{\mathrm{ideal}}\rangle$ and the optimal state $|\psi[\bm{d},\bm{a},\bm{k}_{\mathrm{c}}]\rangle$ are plotted in Fig. 7. The squared overlap between the target and optimal states was found to be $0.992$
Figure 7: Target state $|\psi_{\mathrm{ideal}}\rangle$ for $n_{q}=5$ data qubits and the LC $|\psi[\bm{d},\bm{a},\bm{k}_{\mathrm{c}}]\rangle$ of LF states ($n_{\mathrm{loc}}=3$) that optimally approximates the target state. The contribution from each LF state is also shown.